NOvA running well for the most part. Continuing to take advantage of NuMI downtime to do work. Did have some chiller issues yesterday at Ash River. ArgonCube has been doing well. Continuing to test electronics at LArTF. Module 2 is at LArTF, arrived from CERN without incident so far (haven’t unboxed yet). SENSEI continues to run well. MAGIS making preparations for Laser Transport System vertical test at MINOS. NEXUS is continuing to work on chiller upgrade in the Underground. QUIET continues to prepare for cleanroom build. Will continue to have contractors and employees in for this, let me (Cindy Joe) or Steve Hahn know if questions or access issues. Tom Wicks’ crew is working to remove old MINOS LCW skid today. Thanks to Ben Russell, Joel Fulgham, and rad safety for working with us to help that happen safely. Preparing for epoxy floor installation February 1-4. Tarp has been installed to keep water off the floor. Floor materials are not unsafe but we will likely keep people out of the Underground next weekend to let the fumes dissipate.